OpenStudy (madmerc):

What is the solution to the equation 7^−3x ≈ 9?

OpenStudy (threshsupport):

what does the ^ represent?

OpenStudy (madmerc):

\[7^{-3x} \approx 9\] @threshsupport

OpenStudy (s4sensitiveandshy):

take ln both sides \[ y^x = x \ln(y)\]

OpenStudy (s4sensitiveandshy):

to drop down the power

OpenStudy (madmerc):

i don't understand.. @S4Sensitiveandshy

OpenStudy (s4sensitiveandshy):

\[y^x =z \]to bring the power down I will take log(or ln) both sides \[x \log(y)=\log(z)\] the power rule^^^

OpenStudy (s4sensitiveandshy):

make sense ???

OpenStudy (madmerc):

is it x = 0.295 ? @S4Sensitiveandshy

OpenStudy (s4sensitiveandshy):

hmm how did you get that ??

OpenStudy (arthur_ser):

\[7^{-3x}=9 \rightarrow \log _{7}9=|-3x| \rightarrow \frac{ \ln9 }{ \ln7}=|-3x| \] Solve for x